For a random sample of 12 married couples the age difference (male age minus female age) in years is recorded below. For the below data calculate;
3 0 -4 7 12 2
4 -2 1 20 1 6

For this data calculate the ;
mean
range
interquartile
mean deviation

show all working
1
Expert's answer
2020-04-15T17:56:03-0400

Lets arrange of the values in an ascending order:

"-4,-2,0,1,1,2,3,4,6,7,12,20"


"mean=\bar{x}={1 \over N}\displaystyle\sum_{i=1}^Nx_i="

"={-4-2+0+1+1+2+3+4+6+7+12+20 \over 12}={25 \over 6}\approx4.666667"

"Range=20-(-4)=24"

1st Quartile (Q1): "\dfrac{0+1}{2}=0.5"

2nd Quartile (Q2): "\dfrac{2+3}{2}=2.5"

3rd Quartile (Q3): "\dfrac{6+7}{2}=6.5"

"IQR = Q_3 - Q_1 = 6.5 - 0.5 = 6"


Or

"Q_1=(\dfrac{12+1}{4})^{th}Term=0"

"Q_3=(\dfrac{3(12+1)}{4})^{th}Term=7"


"IQR = Q_3 - Q_1 = 7-0 = 7"


"\begin{matrix}
 x_i & x_i-\bar{x} & |x_i-\bar{x}| \\
 -4 & -49/6 & 49/6 \\
 -2 & -37/6 & 37/6\\
 0 & -25/6 & 25/6\\
 1 & -19/6 & 19/6\\
 1 & -19/6 & 19/6\\
 2 & -13/6 & 13/6\\
 3 & -7/6 & 7/6\\
 4 & -1/6 & 1/6\\
 6 & 11/6 & 11/6\\
 7 & 17/6 & 17/6\\
 12 & 47/6 & 47/6\\
 20 & 95/6 & 95/6

\end{matrix}"

"mean\ deviation={1 \over N}\displaystyle\sum_{i=1}^N|x_i-\bar{x}|="

"={1 \over 12}({340 \over 6})={85 \over 18}\approx4.722222"